Determine the best level of output for a perfectly competitive firm that sells its product at P = $4 and faces TC = 0.04Q3– 0.9Q2 + 10Q + 5. Will the firm produce this level of output? Why?                                                                               2 marks      

1
Expert's answer
2021-01-12T08:00:59-0500

Total revenue= "P*Q" ="4*Q" ="4Q"


"Profit= TR-TC"


"Profit=4Q-(0.04Q^3-0.9Q^2+10Q+5)"


"\frac{dP} {dQ}" ="\u20130.12Q^{2} +1.8Q \u20136"


"\frac{dP} {dQ} =0"


"\u20130.12Q^{2} +1.8Q\u20136""=0" divide across by 0.12


"\frac{-0.12Q^2} {0.12} + \frac{1.8} {0.12}-\frac{6} {0.12} =\frac{0} {0.12}"


"\u2013Q^{2} +15Q-50=0"


"(-Q+10)(Q-5)=0"


"(Q=5) and (Q=10)"


For Q=5; Profit = "\u20130.04(5)^{3} +0.9(5)^{2} \u20136(5)\u20135" = –17.5


For Q=10; Profit = "\u20130.04(10)^{3} +0.9(10)^{2} \u20136(10)\u20135" = –15


No. The firm would not produce this level of output since there are losses when selling the output at $4 dollars per unit.